8. a) A wire loop of resistance 10 and radius 10 cm is kept in the plane of this
paper in a uniform magnetic field . B 
The direction of B 
is perpendicular to the
plane of the page and points out of it and its magnitude is increasing at the rate
of 0.50Ts . 1 Determine the magnitude and direction of the induced current in the
loop.
1
Expert's answer
2020-05-29T09:58:19-0400

Given data

Resistance is "R=10 \Omega"

Rate of magnetic field is "\frac{dB}{dt}=0.50 T/s"

The induced emf in the loop is

"E=-A \frac{dB}{dt}=- \pi r^2 \frac{dB}{dt}"

"=- \pi (0.10 m)^2 (0.50 T/s)"

"=-0.0157 V"

So the magnitude of induced emf is "E=0.0157 V" .

Magnitude of induced current is

"I= \frac{E}{R}=\frac{0.0157 V}{10 \Omega} =1.57*10^{-3} A=1.57 mA"

According to Lenz's law, the direction of induced emf is Clockwise.